OpenStudy (celinegirl):

Simplify square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses. 14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (celinegirl):

I can post a snapshot if its confusing

OpenStudy (anonymous):

yes please post a pic lol

OpenStudy (anonymous):

I think I understand it but I just want to double check

OpenStudy (celinegirl):

Sure (:

OpenStudy (anonymous):

I would say first you need to solve for the square roots. So the square root of 5 is 2.2 rounded to the nearest tenth and the square root of 2 is 1.4 rounded to the nearest tenth.

OpenStudy (anonymous):

Now you can rewrite your expression as \[2.2(10-4(1.4))\]

OpenStudy (calculusxy):

\[\sqrt{5}(10 - 4\sqrt{2})\] \[10\sqrt{5} - 4\sqrt{10}\]

OpenStudy (anonymous):

now we solve using pemdas. So starting the parentheses with the multiplication we have 4*1.4=5.6. Now we solve for the rest of the parentheses with 10-5.6=4.4.

OpenStudy (anonymous):

Now all we have to do is multiply 2.2*4.4=9.68... So does that match up with any of your options?

OpenStudy (anonymous):

Hopefully I am right.... Sorry if I am not

OpenStudy (celinegirl):

@kimbllse777 C right?

OpenStudy (shamil98):

no. the answer would be as calculusxy posted since it's not an option go with none of the above.

OpenStudy (celinegirl):

I just dont want to get it wrong

OpenStudy (shamil98):

it's the last option, D. lol.

OpenStudy (celinegirl):

@shamil98 ok thank you